Can you please check my work and let me know if I did this right?
PCB 3063L ProBof or aAs 1:2:1 10. In humans, the rh factor is inherited as a dominant gene (R). Individuals with this llele are referred to as rh positive. Two (2) heterozygous individuals (Rr) decided to reproduce. What is the probability of four (4) rh positivechildrenzu- a. b. What is the probability of four (4) rh negative boys? [1] nuy prop r 0.02% nue rc. What is the probability of two (2) rh positive boys and two (2) rh negative girls [] oro . 22% d. If they have five (5) children, what is the probability that the first two (2) will be rh negative and the last three (3) will be rh positive? [11 posifvè: 34 2.6%. e. What is the probability of three (3) rh positive children out of eight (8) children? [11 You will need to use the binomial expansion formula for this problem: - s as 31(R-3) I0320x 042Vas x.DOb4ausbas 220 or275%
Answer #1

rh positive (R) > rh negative (r)
rh positive genotypes = RR, Rr, rR
rh negative genotypes = rr

Genotypes of the parents:
Heterozygous rh positive = Rr

As per Mendelian Laws of Inheritance (Segregation, Independent Assortment, and Dominance):
Cross There are 2 distinct phenotypes rR 75 25 % Rr
Probability of an rh positive child = 3/4
Probability of an rh negative child = 1/4
Probability of a boy = Probability of a girl = 1/2

ANSWER 10.
Answer a.

P of 4 rh+ children = (3/4)4 = 31.64%

Answer b.
P of 4 rh- boys = ((1/4) * (1/2))4 = 0.0244%

Answer c.
P of 2 rh+ boys & 2 rh- girls = 4C2? (3/4 * 1/2)2 * (1/4 * 1/2)2 = 1.318% The order in which the kids are born also matters hence the 4C2

Answer d.
P of first 2 to be rh- & last 3 rh+ = (1/4)2 * (3/4)3 = 2.63%

Answer e.
P of 3 rh+ children out of 8 children = ?8C3? (3/4)3 * (1/4)5 = 2.307%